# Zero Trust Architecture: A Systematic Literature Review

### v1

@@ 5.1.2 Communication Networks
The introduction of 6G technology after 5G, wireless communication has entered an exciting new era that will revolutionise connectivity like never before [64]. This technology has the potential to completely alter the limits of communication, with predicted data speeds of up to ten terabits per second and low latency [97]. However, as these networks become more complex and linked, they might be vulnerable to new cybersecurity threats [98]. Hence, Implementing extra security measures like encryption, continuous monitoring, and multi-factor authentication (MFA), which are key components of ZTA, is essential to addressing the security threats posed by these networks [99].
@@
Various studies have shown the applicability of ZTA in fortifying the security defences of 5G/6G enabled networks. For instance, Nurun Nahar et al. [14] conducted a survey that examined the limitations of traditional perimeter-based security models in securing 6G networks, highlighting their challenges and vulnerabilities in the face of sophisticated cyber threats. Accordingly, the authors indicated the potential and applicability of ZTA as a robust security framework that can address internal threats while enhancing the security status of 6G enabled networks. Similarly, the work presented in [100] further underscore the need for integrating ZTA into 5G/6G communication networks while highlighting the limitations of PBSM. The authors proposed an intelligent ZTA framework which is integrated with AI engines to dynamically assess and authorize access request based on various factors such security policies, and subject privileges. Key components of the proposed framework include an AI trust algorithm for access authorization, a risk assessment engine utilizing graph neural networks for security state evaluation, and a user AI engine that analyzes network traffic to maintain a high confidence level in user access.
@@
In addition, a zero-trust authentication system that does not rely on the presence of secure channels was developed by Seunghwan et al. [64] for 6G enabled IoT environments. Using a lightweight attribute-based encryption (ABE) system for dynamic access control, the framework allows the control of access rights based on the status of user equipment (UE) and leverages blockchain technology to enable mutual authentication between network elements, guaranteeing trustworthiness in a decentralised manner. The design addresses the particular difficulties brought about by the dispersed and varied nature of 6G-enabled IoT environments, with the goal of improving security and efficiency in 6G networks by providing robust authentication methods and fine-grained access control. Xu et al. [101] proposed a software-defined ZTA for 6G networks which adaptively manages access control across multiple control domains. A blockchain-based third-party security services to enhance resource management and support decentralized identity management was used. The effectiveness and robustness of the proposed ZTA framework are validated through simulations, demonstrating its capability to counteract security threats in the evolving landscape of 6G networks.
@@
Satellite communication systems support critical operations, including military communication, remote sensing, and global navigation. Their reliance on long-range, often public, channels makes them vulnerable to eavesdropping, signal spoofing, and jamming attacks [102]. ZTA can enhance the security of satellite communication by enforcing encrypted data exchanges, mutual authentication, and other secrurity features between ground stations, satellites, and end-user devices. For example, in [103], Shiva proposed a ZTA framework for securing Low Earth Orbit (LEO) satellite communication networks called orbital ZTA. The framework was developed based on a directive in collaboration with commercial satellite communication operators. It follows the NIST ZTA framework but decouples the control plane, which were trained on-board on LEO satellites using machine learning, from the data plane modules, which were trained locally within each satellite. In addition, the framework incorporates an orbital policy engine (OPE) that evaluates access decisions based on enterprise policy, enabling fine-grained access control and continuous trust evaluation, thereby enhancing the overall security posture of satellite communications against emerging threats. Furthermore, a ZTA framework for securing satellite networks that uses edge intelligence for continuous authentication is proposed in [72]. According to the authors, the framework considers environment, behavior, and physical entity of the satellite networks in addition to the subject and object commonly focused on by traditional ZTA models. Also, the framework ensures proactive and continuous authentication through periodic monitoring and re-evaluation of variable attributes, and a Neural-Backed Decision Trees (NBDTs) based intelligence algorithm designed to enhance the accuracy of authentication. Through a simulation testbed, the framework is shown to demonstrate significant authentication accuracy increment and has out-performed the conventional attribute-based access control (ABAC) model.
@@
Similarly, to enable distributed and secure authentication in a network environment, Javier et al. [107] proposed a distributed, and privacy-preserving multi-factor authentication (MFA) framework designed for zero trust networks based on blockchain.To improve system resilience and remove single points of failure, the framework makes use of a Distributed Authentication Mechanism (DAM), in which every node in the blockchain network participates in the authentication process. To ensure privacy and confidentiality, the framework uses a partial secret generation approach to generate an OTP and Zero-Knowledge Proofs (ZKP) to validate the OTP without disclosing its value. Furthermore, it employs Elliptic Curve Cryptography (ECC) and Elliptic Curve Digital Signature Algorithm (ECDSA) to protect communications and employs modified ERC721 tokens as distinct, non-transferable authentication tokens with a predetermined validity time. The authors indicate the effectiveness of the framework through experimental results and comparison with existing works.
@@
Open Radio Access Network (O-RAN) is a revolutionary approach to building mobile networks that prioritizes openness, interoperability, and flexibility [108]. Unlike traditional RAN architectures that relied on proprietary, and single-vendor solutions, O-RAN embraces open interfaces and standardized protocols which allows operators to mix and match components from different vendors [70]. However, this increased flexibility and openness also introduce new security challenges such as supply chain complexity due to multiple vendors, and increased attack surface [71]. To address these challenges, ZTA has been proposed to enhance the security within the O-RAN framework.
In [108], a blockchain-based ZTA is proposed to enhance the trustworthiness and security in O-RAN which enables the authentication and verification of applications managed by O-RAN vendors to prevent unauthorized access from malicious applications. The authors demonstrated the effectiveness of the framework on both public and private blockchain environment. Similarly, in [100], while acknowledging the security vulnerabilities of O-RAN, a framework based on some key ZT principles such as continuous verification, least privilege, and data-centric security, and three intelligent components, ”Intelligent Policy Engine (IPE)”, ”Intelligent Network Security State Analysis (INSSA)”, and ”Intelligent Agent/Portal (IGP)”, are used for dynamic approach to securing O-RAN including ”Continuous Multi-Factor Authentication (CMFA)” to enhance trust verification within the network. The framework models real-time access authorization decisions, security state of the network, as well as network awareness by analyzing network traffic patterns.
Furthermore, Hajar et al. [70], [71] leverages two deep reinforcement learning techniques in the two articles - Q-learning and Deep Sarsa, respectively, to train and evaluate their models on UNSW dataset in their proposed ZTA frameworks for O-RAN security. Both frameworks have been shown to be effective in securing the O-RAN network using some performance metrics. Similarly, in [109], Transformers were used in conjunction with key ZT principles to develop an anomaly detection model for analysis of real-time time series 5G-NIDD dataset to enhance the security of O-RAN networks. The model is shown to have superior performance compared to other models.
@@ 5.1.3 Finance and Commerce
There is no doubt that digital technologies have developed at an unprecedented pace in recent years and the finance industry is among the most impacted industries [110]. Consequently, as the financial institutions navigate through the realm of digital transformation, their attack surface and susceptibility to an increasingly complex cyber threat landscape is substantial and this creates an urgent need for a robust and dynamic security architecture [73]. More recently in 2023, organizations targetted during the ”MOVEit Vulnerability Exploitation” included financial institutions. In May, a ransomware group utilized vulnerabilities in Progress Software’s MOVEit Transfer tool that led to the compromise of sensitive data from a large number of companies around the globe such as banks and payment processing companies [111].
In order to offer some solutions regarding the risks that the financial sector encounters, Clement et al. [73] has put forward a ZTA Framework which features Blockchain technology and robust Identity and Access Management (IAM) processes to ensure secure and immutable user authentication and authorization. They incorporate comprehensive threat detection through various security components, adopts an adaptive security posture to respond to evolving threats, and aims for operational efficiency by maintaining high throughput and low latency. Overall, it offers an extensive defensive structure formulated to effectively counter cyber threats while adhering to compliance regulations such as Anti-Money Laundering (AML) and Know Your Customer (KYC).. In a separate article, the authors proposed another framework based on Blockchain, evaluated by OMNeT++ simulations and improved through Ethereum-Ganache [112]. Similarly, in another article by the authors [113], they proposed a ZTA tailored for the financial sector which aims to address three key security areas - IAM, device and network security, and data protection, through the implementation of various security concepts such as MFA, IDS, data loss prevention (DLP), network segmentation, e.t.c. The testing results from a demo bank app show that the framework effectively addresses the identified security challenges in all the three areas. However, challenges remain, such as integrating the framework with legacy systems and balancing security with user experience.
@@ 5.1.4 Internet of Things
The internet of things (IoT) has being a key enabler and driving force for today’s digital transformation and industrial revolution [117]. Current networks have grown as a result of the IoT widespread acceptance [118], with an increasing number of terminals using resources, exchanging information, and conducting data transactions both inside and outside of network perimeters [119]. Coupled with the surge in cyber attacks, this has led the traditional PBSM less significant [79]. Various solutions have been proposed by researchers over the years to mitigate the security issues and challenges of IoT given their unique features and diverse characteristics, and more recently, the focus has being on ZTA.
@@
Furthermore, other researchers explore how ZT principles can be implemented in IoT and Industrial IoT environments by investigating secure access control technologies, and evaluating emerging authentication methods. More specifically, Yanliu [120] conduct a research that analyzes access control methods used in ZT models and performed an empirical experiments to assess their feasibility in real industrial IoT. Chanapha and Kuo-Hui [117] discuss on the various authentication methods used in ZTA for ensuring IoT security, highlight some strategies for effective implementation, and introduced the performance evaluation criteria necessary for assessing the effectiveness of authentication methods within IoT environments. Moreover, in [121], Theo et al proposes a ZTA framework for trust-aware continuous authorization to secure IoT environment such as smart homes. The framework leverages a message-centric architecture based on a publish/subscribe model and integrates a trust evaluation engine into ABAC authorization policy engine. It ensures contextualized and secure session management through continuous monitoring and re-evaluation of both trust levels and authorization policies which are modified whenever changes to features about subjects, resources, environment, or the trust levels are detected throughout the authorization period. The performance of a prototype of the framework was validated under three test conditions mostly in terms of the time taken to receive re-evaluation against number of attributes. Lastly, on another different perspective, Chunwen et al. [45] conduct an elaborate bibliometric analysis to the assess the research landscape of ZTA for IoT security. The authors explored various vulnerabilities associated with IoT devices and the potential of ZT strategy to mitigate those security concerns.
Additionally, Jose et al [122] extended application of ZTA to computing continuum, a progressive convergence of cloud computing and IoT. In [122], they emphasized on decentralized identity management for exchanging resources in a computing continuum through robust authentication and authorization mechanisms. Using Self-Sovereign Identity (SSI) principles, individuals retain ownership and control of their digital identities in the heterogeneous environment, ensuring privacy and trust management. The proposed system employs blockchain for decentralization, fostering trust and interoperability across domains, and privacy-preserving techniques such as Zero Knowledge Proofs that allow selective disclosure of sensitive information. The authors implemented several modules and validated the proposed framework according to FLUIDOS project scope.
@@ 5.1.5 Online Communities
From the reviewed articles, A ZT Access Control (ZTAC) framework was proposed by Taisho et al. [63] in order to improve user authenticity verification in teleworking settings. In order to continuously evaluate user behaviour and confirm the legitimacy of account operators, the framework integrates behavioural and cognitive web-biometrics, such as tracking click actions and keyboard typing speed. The authors assessed the proposed framework in terms of attack patterns and system performance after defining four remote access strategies that attackers frequently employ. Similarly, Shiva et al. [30] introduced a framework based on Blockchain and federated learning that dynamically adapts to changing user situations with realible trust computation to improve remote work and collaboration. Some Clustering techniques and incremental anomaly detection were to distinguish and share local and global anomalies among nodes. Also, in [125], a ZTA framework is proposed to secure remote access to online resources using Zero Trust Network Access (ZTNA), specifically using Twingate, and comprises of several components including MFA, encryption, and device security policies. While considering an online laboratory as a case study, the authors highlight the efficacy of the approach over traditional VPNs in establishing least privilege principle, providing secure access, and mitigating various VPN-related vulnerabilities.
Furthermore, the concept of ZTA is extended to social media platforms. As shown in [83], Jiawei et al. proposed a framework that ensure the security of user identity and the authenticity of the shared content using technologies such as OAuth 2.0 and OpenID Connect for secure user authentication, biometric recognition for enhanced identity verification, and digital signature technology to verify content integrity. The framework adopts a ZT security model, which emphasizes continuous verification of user identities and access requests, and integrates deep learning techniques to manage and analyze multi-modal content efficiently. Another application of ZTA is in secure reviewing of articles by reviewers. In [60], Pooja et al. proposed an access control framework for secure reviewing and sharing of data which integrates ZTA principles to enhance confidentiality in scientific collaborations so that a manuscript is only accessible to the assigned reviewer. The framework uses blockchain technology to facilitate secure and transparent data sharing among the concerned parties which consists of the author, reviewer, editor, and publisher and incorporates advanced encryption methods, such as AES encryption to protect sensitive data during transit. In addition, the framework also includes mechanisms for author feedback on review comments, which adds value to the peer review process and aids in the selection of future reviewers.

### v2

@@ 5.1.2 Communication Networks
The introduction of 6G technology after 5G, wireless communication has entered an exciting new era that will revolutionize connectivity like never before [70]. This technology has the potential to completely alter the limits of communication, with predicted data speeds of up to ten terabits per second and low latency [103]. However, as these networks become more complex and linked, they might be vulnerable to new cybersecurity threats [104]. Hence, Implementing extra security measures like encryption, continuous monitoring, and multifactor authentication (MFA), which are key components of ZTA, is essential to addressing the security threats posed by these networks [105].
@@
Various studies have shown the applicability of ZTA in fortifying the security defenses of 5G/6G enabled networks. For instance, Nurun Nahar et al. [14] conducted a survey that examined the limitations of traditional perimeter-based security models in securing 6G networks, highlighting their challenges and vulnerabilities in the face of sophisticated cyber threats. Accordingly, the authors indicated the potential and applicability of ZTA as a robust security framework that can address internal threats while enhancing the security status of 6G enabled networks. Similarly, the work presented in [106] further underscores the need for integrating ZTA into 5G/6G communication networks while highlighting the limitations of PBSM. The authors proposed an intelligent ZTA framework which is integrated with AI engines to dynamically assess and authorize access request based on various factors such security policies, and subject privileges. Key components of the proposed framework include an AI trust algorithm for access authorization, a risk assessment engine utilizing graph neural networks for security state evaluation, and a user AI engine that analyzes network traffic to maintain a high confidence level in user access.
@@
In addition, a zero-trust authentication system that does not rely on the presence of secure channels was developed by Seunghwan et al. [70] for 6G enabled IoT environments. Using a lightweight attribute-based encryption (ABE) system for dynamic access control, the framework allows the control of access rights based on the status of user equipment (UE) and leverages blockchain technology to enable mutual authentication between network elements, guaranteeing trustworthiness in a decentralized manner. The design addresses the particular difficulties brought about by the dispersed and varied nature of 6G-enabled IoT environments, with the goal of improving security and efficiency in 6G networks by providing robust authentication methods and fine-grained access control. Xu et al. [107] proposed a software-defined ZTA for 6G networks which adaptively manages access control across multiple control domains. A blockchain-based third-party security services to enhance resource management and support decentralized identity management was used. The effectiveness and robustness of the proposed ZTA framework are validated through simulations, demonstrating its capability to counteract security threats in the evolving landscape of 6G networks.
@@
Satellite communication systems support critical operations, including military communication, remote sensing, and global navigation. Their reliance on long-range, often public, channels makes them vulnerable to eavesdropping, signal spoofing, and jamming attacks [108]. ZTA can enhance the security of satellite communication by enforcing encrypted data exchanges, mutual authentication, and other security features between ground stations, satellites, and end-user devices. For example, in [109], Shiva proposed a ZTA framework for securing Low Earth Orbit (LEO) satellite communication networks called orbital ZTA. The framework was developed based on a directive in collaboration with commercial satellite communication operators. It follows the NIST ZTA framework but decouples the control plane, which were trained on-board on LEO satellites using machine learning, from the data plane modules, which were trained locally within each satellite. In addition, the framework incorporates an orbital policy engine (OPE) that evaluates access decisions based on enterprise policy, enabling fine-grained access control and continuous trust evaluation, thereby enhancing the overall security posture of satellite communications against emerging threats. Furthermore, a ZTA framework for securing satellite networks that uses edge intelligence for continuous authentication is proposed in [78]. According to the authors, the framework considers environment, behavior, and physical entity of the satellite networks in addition to the subject and object commonly focused on by traditional ZTA models. Also, the framework ensures proactive and continuous authentication through periodic monitoring and re-evaluation of variable attributes, and a Neural-Backed Decision Trees (NBDTs) based intelligence algorithm designed to enhance the accuracy of authentication. Through a simulation testbed, the framework is shown to demonstrate significant authentication accuracy increment and has outperformed the conventional attribute-based access control (ABAC) model.
@@
Similarly, to enable distributed and secure authentication in a network environment, Javier et al. [113] proposed a distributed, and privacy-preserving multifactor authentication (MFA) framework designed for zero trust networks based on blockchain. To improve system resilience and remove single points of failure, the framework makes use of a Distributed Authentication Mechanism (DAM), in which every node in the blockchain network participates in the authentication process. To ensure privacy and confidentiality, the framework uses a partial secret generation approach to generate an OTP and Zero-Knowledge Proofs (ZKP) to validate the OTP without disclosing its value. Furthermore, it employs Elliptic Curve Cryptography (ECC) and Elliptic Curve Digital Signature Algorithm (ECDSA) to protect communications and employs modified ERC721 tokens as distinct, non-transferable authentication tokens with a predetermined validity time. The authors indicate the effectiveness of the framework through experimental results and comparison with existing works.
@@
Open Radio Access Network (O-RAN) is a revolutionary approach to building mobile networks that prioritizes openness, interoperability, and flexibility [114]. Unlike traditional RAN architectures that relied on proprietary, and single-vendor solutions, O-RAN embraces open interfaces and standardized protocols which allow operators to mix and match components from different vendors [76]. However, this increased flexibility and openness also introduce new security challenges such as supply chain complexity due to multiple vendors, and increased attack surface [77]. To address these challenges, ZTA has been proposed to enhance the security within the O-RAN framework.
In [114], a blockchain-based ZTA is proposed to enhance trustworthiness and security in O-RAN which enables the authentication and verification of applications managed by O-RAN vendors to prevent unauthorized access from malicious applications. The authors demonstrated the effectiveness of the framework on both public and private blockchain environments. Similarly, in [106], while acknowledging the security vulnerabilities of O-RAN, a framework based on some key ZT principles such as continuous verification, least privilege, and data-centric security, and three intelligent components, ”Intelligent Policy Engine (IPE)”, ”Intelligent Network Security State Analysis (INSSA)”, and ”Intelligent Agent/Portal (IGP)”, are used for dynamic approach to securing O-RAN including ”Continuous multifactor Authentication (CMFA)” to enhance trust verification within the network. The framework models real-time access authorization decisions, security state of the network, as well as network awareness by analyzing network traffic patterns.
Furthermore, Hajar et al. [76], [77] leverage two deep reinforcement learning techniques in the two articles - Q-learning and Deep Sarsa, respectively, to train and evaluate their models on UNSW dataset in their proposed ZTA frameworks for O-RAN security. Both frameworks have been shown to be effective in securing the O-RAN network using some performance metrics. Similarly, in [115], Transformers were used in conjunction with key ZT principles to develop an anomaly detection model for the analysis of real-time time series 5G-NIDD dataset to enhance the security of O-RAN networks. The model is shown to have superior performance compared to other models.
@@ 5.1.3 Finance and Commerce
There is no doubt that digital technologies have developed at an unprecedented pace in recent years and the finance industry is among the most impacted industries [116]. Consequently, as the financial institutions navigate through the realm of digital transformation, their attack surface and susceptibility to an increasingly complex cyber threat landscape is substantial, and this creates an urgent need for a robust and dynamic security architecture [79]. More recently in 2023, organizations targeted during the ”MOVEit Vulnerability Exploitation” included financial institutions. In May, a ransomware group utilized vulnerabilities in Progress Software’s MOVEit Transfer tool that led to the compromise of sensitive data from a large number of companies around the globe such as banks and payment processing companies [117].
In order to offer some solutions regarding the risks that the financial sector encounters, Clement et al. [79] has put forward a ZTA Framework which features Blockchain technology and robust Identity and Access Management (IAM) processes to ensure secure and immutable user authentication and authorization. They incorporate comprehensive threat detection through various security components, adopts an adaptive security posture to respond to evolving threats, and aims for operational efficiency by maintaining high throughput and low latency. Overall, it offers an extensive defensive structure formulated to effectively counter cyber threats while adhering to compliance regulations such as Anti-Money Laundering (AML) and Know Your Customer (KYC). In a separate article, the authors proposed another framework based on Blockchain, evaluated by OMNeT++ simulations and improved through Ethereum-Ganache [118]. Similarly, in another article by the authors [119], they proposed a ZTA tailored for the financial sector which aims to address three key security areas - IAM, device and network security, and data protection, through the implementation of various security concepts such as MFA, IDS, data loss prevention (DLP), network segmentation, etc. The testing results from a demo bank app show that the framework effectively addresses the identified security challenges in all the three areas. However, challenges remain, such as integrating the framework with legacy systems and balancing security with user experience.
@@ 5.1.4 Internet of Things
The internet of things (IoT) has been a key enabler and driving force for today’s digital transformation and industrial revolution [123]. Current networks have grown as a result of the IoT widespread acceptance [124], with an increasing number of terminals using resources, exchanging information, and conducting data transactions both inside and outside of network perimeters [125]. Coupled with the surge in cyber attacks, this has led the traditional PBSM less significant [85]. Various solutions have been proposed by researchers over the years to mitigate the security issues and challenges of IoT given their unique features and diverse characteristics, and more recently, the focus has been on ZTA.
@@
Furthermore, other researchers explore how ZT principles can be implemented in IoT and Industrial IoT environments by investigating secure access control technologies, and evaluating emerging authentication methods. More specifically, Yanliu [126] conduct a research that analyzes access control methods used in ZT models and performed an empirical experiments to assess their feasibility in real industrial IoT. Chanapha and Kuo-Hui [123] discuss on the various authentication methods used in ZTA for ensuring IoT security, highlight some strategies for effective implementation, and introduced the performance evaluation criteria necessary for assessing the effectiveness of authentication methods within IoT environments. Moreover, in [127], Theo et al propose a ZTA framework for trust-aware continuous authorization to secure IoT environment such as smart homes. The framework leverages a message-centric architecture based on a publish/subscribe model and integrates a trust evaluation engine into ABAC authorization policy engine. It ensures contextualized and secure session management through continuous monitoring and re-evaluation of both trust levels and authorization policies which are modified whenever changes to features about subjects, resources, environment, or the trust levels are detected throughout the authorization period. The performance of a prototype of the framework was validated under three test conditions mostly in terms of the time taken to receive re-evaluation against number of attributes. Lastly, on another different perspective, Chunwen et al. [45] conduct an elaborate bibliometric analysis to the assess the research landscape of ZTA for IoT security. The authors explored various vulnerabilities associated with IoT devices and the potential of ZT strategy to mitigate those security concerns.
Additionally, José et al [128] extended application of ZTA to computing continuum, a progressive convergence of cloud computing and IoT. In [128], they emphasized on decentralized identity management for exchanging resources in a computing continuum through robust authentication and authorization mechanisms. Using Self-Sovereign Identity (SSI) principles, individuals retain ownership and control of their digital identities in the heterogeneous environment, ensuring privacy and trust management. The proposed system employs blockchain for decentralization, fostering trust and interoperability across domains, and privacy-preserving techniques such as Zero Knowledge Proofs that allow selective disclosure of sensitive information. The authors implemented several modules and validated the proposed framework according to FLUIDOS project scope.
@@ 5.1.5 Online Communities
From the reviewed articles, A ZT Access Control (ZTAC) framework was proposed by Taisho et al. [69] in order to improve user authenticity verification in teleworking settings. In order to continuously evaluate user behaviour and confirm the legitimacy of account operators, the framework integrates behavioural and cognitive web-biometrics, such as tracking click actions and keyboard typing speed. The authors assessed the proposed framework in terms of attack patterns and system performance after defining four remote access strategies that attackers frequently employ. Similarly, Shiva et al. [30] introduced a framework based on Blockchain and federated learning that dynamically adapts to changing user situations with reliable trust computation to improve remote work and collaboration. Some clustering techniques and incremental anomaly detection were to distinguish and share local and global anomalies among nodes. Also, in [131], a ZTA framework is proposed to secure remote access to online resources using Zero Trust Network Access (ZTNA), specifically using Twingate, and comprises of several components including MFA, encryption, and device security policies. While considering an online laboratory as a case study, the authors highlight the efficacy of the approach over traditional VPNs in establishing least privilege principle, providing secure access, and mitigating various VPN-related vulnerabilities.
Furthermore, the concept of ZTA is extended to social media platforms. As shown in [89], Jiawei et al. proposed a framework that ensures the security of user identity and the authenticity of the shared content using technologies such as OAuth 2.0 and OpenID Connect for secure user authentication, biometric recognition for enhanced identity verification, and digital signature technology to verify content integrity. The framework adopts a ZT security model, which emphasizes continuous verification of user identities and access requests, and integrates deep learning techniques to manage and analyze multi-modal content efficiently. Another application of ZTA is in secure reviewing of articles by reviewers. In [66], Pooja et al. proposed an access control framework for secure reviewing and sharing of data which integrates ZTA principles to enhance confidentiality in scientific collaborations so that a manuscript is only accessible to the assigned reviewer. The framework uses blockchain technology to facilitate secure and transparent data sharing among the concerned parties which consists of the author, reviewer, editor, and publisher and incorporates advanced encryption methods, such as AES encryption to protect sensitive data during transit. In addition, the framework also includes mechanisms for author feedback on review comments, which adds value to the peer review process and aids in the selection of future reviewers.